# Linear Combination of Saved Checkpoints Makes Consistency and Diffusion Models Better

### v2

@@ Appendix 0.A Additional Examples of the Metric Landscape
Appendix 0.B Experimental Details and Additional Results
0.B.1 Experimental Details
0.B.1.1 Training Setting
@@
Since LCSC employs weight checkpoints, we endeavored to replicate the training processes of the baseline models. For DMs, we were able to reproduce the results reported in the original papers successfully. However, for CMs, our training outcomes on CIFAR10 were slightly inferior to those documented in the original papers. Additionally, due to resource constraints, we opted for a smaller batch size than the original configuration when training on ImageNet. In the results tables, we present both our training outcomes and the results reported in the original papers, with the latter indicated as released. The specifics of the experimental setups are detailed below.
@@
For the experiments with DM, we utilize the DDIM [37] codebase (https://github.com/ermongroup/ddim) on CIFAR-10, adhering to the default configuration settings. For ImageNet-64, we employ the iDDPM [28] codebase (https://github.com/openai/improved-diffusion), setting the batch size to 512, the noise schedule to cosine, and maintaining other hyperparameters at their default values. During sampling, the DPM-Solver [24] (https://github.com/LuChengTHU/dpm-solver) is applied, conducting 15 timesteps of denoising using the default configuration for the respective setting. The reproduced results are on par with or slightly surpass those achieved using Euler integration, as reported in the original papers.
@@
For the experiments involving CM, we adhere closely to the configurations detailed in Table 3 of the original CM paper [39]. On CIFAR-10, instead of using the NCSNPP model [40] from CM’s official implementation, we opt for the EDM [15] architecture. This decision stems from the official CM code for CIFAR-10 being implemented in JAX while our experiments are conducted using PyTorch. For further details, refer to https://github.com/openai/consistency_models_cifar10. This difference in implementation may partly explain the discrepancies between our replication results and those reported in Table 1 of CM’s original paper [39]. On ImageNet-64, we follow the official implementation, employing the ADM [4] architecture. Due to resource constraints, we train the model with a smaller batch size of 256 instead of 2048 as used in the original study, resulting in worse outcomes. However, as indicated in Tab. 2, applying LCSC to models trained with a reduced batch size can still achieve performance comparable to models trained with a larger batch size.
0.B.1.2 Search Setting
@@
At each selected training iteration, historical weights are leveraged within designated window sizes and intervals. For CM, checkpoints have a window size of 40K with an interval of 100 on CIFAR-10, and a window size of 20K with an interval of 100 on ImageNet-64. DM is assigned a window size of 50K and an interval of 200 for both datasets. FID calculation for DM on ImageNet-64 utilizes a sample of 5K images, while 10K images are used for all other configurations. An evolutionary search spanning 2K iterations is applied consistently across all experimental setups.
0.B.2 Search Cost Estimation
@@
Next, we examine the overall training and search time costs as outlined in Tab. 6, and we delve into the speedup ratio detailed in Tabs. 7 and 8. Given that DMs necessitate multiple timesteps for sampling, the overall search cost for DM proves to be non-negligible when compared to the training cost. This results in no observable speedup in convergence upon applying LCSC. However, this search cost is considered manageable and warranted for the anticipated improvements to the final model convergence.
In contrast, for CMs, the search cost is substantially lower than the training cost, which often results in a markedly faster convergence when combining training with LCSC than with training alone. The precise speedup ratios for CM training are detailed in Tabs. 7 and 8. It should be noted that the figures in these tables may exhibit slight discrepancies from those reported in Sec. 5 due to the latter being profiled in an environment with more variables and potential fluctuations.
@@
0.B.3 Hyperparameter Study
In Sec. 5.3, we investigate the sensitivity of several hyperparameters, including the number of search iterations, the number of samples, and the interval between two weight checkpoints $(n_{k}-n_{k-1})$. This section focuses on the impact of window size $(n_{K}-n_{1})$ on the performance of our search algorithm. Given that DM operates iteratively during sampling and a larger NFE (i.e. the number of timesteps for the ODE solver) typically leads to higher generation quality, we also examine the effects of varying NFE during the search within the DM context. The findings are detailed in Table 9.
The results reveal that, despite earlier models being further from convergence, a sufficiently large window for accessing historical weights proves advantageous. Furthermore, conducting a lower NFE during the search in the DM context results in a similar reduction in FID but a smaller improvement in IS. This suggests that the search output’s generality across different metrics diminishes when the generated samples during the search are less accurate, i.e., exhibit significant truncation error.
@@ 0.B.3 Hyperparameter Study
0.B.4 Detailed Results of EMA Rate Grid Search
For the final training model of each experimental configuration, we conduct a comprehensive sweep across a broad range of EMA rates, presenting the optimal outcomes in Sec. 5. The exhaustive results are detailed in Tab. 10. In the majority of scenarios, the default EMA rate employed in the official implementations of our baseline models [39, 8, 28] yields slightly inferior performance compared to the best EMA rate identified. Nevertheless, the findings demonstrate that LCSC consistently surpasses all explored EMA rates, underscoring the limitations of EMA as a strategy for weight averaging.
@@ 0.B.4 Detailed Results of EMA Rate Grid Search
0.B.5 Evaluation with FID@CLIP
In this section, we present the results of FID@CLIP (i.e., FID computed within the feature space of a CLIP model) to demonstrate that LCSC facilitates improvements beyond the inception feature space. We utilize the visual component of a CLIP model based on ViT-B/32. Detailed outcomes are documented in Tabs. 11 and 12. The findings clearly indicate that LCSC consistently surpasses the performance of the EMA, as measured by FID@CLIP. Notably, models trained for fewer iterations or with smaller batch sizes experience enhancements through LCSC, achieving better performance than the fully converged EMA model as gauged by FID@CLIP.
Table 11: FID on CIFAR10 calculated within the visual feature space of a CLIP model, denoted as FID@CLIP.
Model Method Training Iter Batch Size NFE FID($\downarrow$) FID@CLIP($\downarrow$)
CD EMA 50K 512 1 4.84 28.1
250K 512 1 4.08 20.3
800K 512 1 3.66 17.7
800K 512 2 2.90 13.3
850K 512 1 3.65 17.3
850K 512 2 2.89 13.1
LCSC 50K 512 1 3.18 21.4
250K 512 1 2.76 14.3
800K 512 1 2.42 14.8
800+40K 512 1 2.38 14.1
CT EMA 400K 512 1 12.08 43.0
800K 512 1 9.87 35.8
LCSC 400K 512 1 8.60 29.5
450K 128 1 8.33 25.5
DM EMA 150K 128 15 6.28 46.4
350K 128 15 4.45 41.9
800K 128 15 4.16 42.7
LCSC 150K 128 15 4.76 43.1
350K 128 15 3.56 36.1
800K 128 15 3.18 36.4
Table 12: FID on ImageNet-64 calculated within the visual feature space of a CLIP model, denoted as FID@CLIP.
Model Method Training Iter Batch Size NFE FID($\downarrow$) FID@CLIP($\downarrow$)
CD EMA 300K 256 1 7.70 30.0
600K 256 1 7.30 29.1
650K 256 1 7.17 28.4
(released) 600K 2048 1 6.31 25.5
LCSC 300K 256 1 5.71 28.2
600+20K 256 1 5.54 24.9
CT EMA 600K 256 1 16.6 49.8
1000K 256 1 15.6 47.5
(released) 800K 2048 1 13.1 47.5
LCSC 600K 256 1 12.1 38.3
DM EMA 250K 512 15 20.8 46.6
500K 512 15 19.8 44.2
LCSC 250K 512 15 16.3 44.8
500K 512 15 15.3 41.5
Appendix 0.C More Insights and Analysis
0.C.1 Analysis of Search Patterns
0.C.1.1 More Examples of Search Patterns
@@
We demonstrate search results on ImageNet-64 in Fig. 7, which share similar patterns with results on CIFAR-10 (refer to Fig. 5). Our analysis in Sec. 6 also holds in these cases.
@@
0.C.1.2 Consistency of Search Results
We evaluate the consistency of search patterns by comparing results across different random seeds. As indicated in Tab. 13, both seeds produce similar outcomes in terms of FID and IS metrics. Yet, Fig. 8 reveals that each seed’s dominant coefficients map to distinct subsets of weight checkpoints, suggesting the presence of multiple high-quality basins within the weight checkpoint subspace. This diversity underscores LCSC’s ability to identify and enhance model performance within one of these basins through a limited number of search iterations (e.g., 2K in our experiments), demonstrating its efficiency and effectiveness.
Figure 8: Visualization of linear combination coefficients obtained by LCSC with different random seeds.
Table 13: Performance of LCSC with different random seeds.
Model Method Training Iter Batch Size NFE Seed FID($\downarrow$) IS($\uparrow$)
CD LCSC 250K 512 1 1 2.76 9.71
2 2.69 9.69
0.C.1.3 The Critical Role of Search Space Configuration
As explored in Sec. 6, restricting averaged weights to the convex hull of candidate checkpoints, where all combination coefficients are non-negative and their sum equals 1, might limit search efficacy. To test this hypothesis, we perform searches under convex combination conditions by clipping coefficients to non-negative values and normalizing their sum to 1. Table 14 presents these findings, with “w/o” indicating no restriction on coefficients (allowing values below zero) and “w” representing the convex hull restriction. The results clearly show superior performance without the convex restriction, underscoring its limiting effect on search outcomes. Furthermore, we discover that normalizing the coefficient sum to 1 is vital for effective search exploration, a practice we continue even after lifting the convex condition, as detailed in Eq. 8.

### v3

@@ Appendix D Additional Examples of the Metric Landscape
Since we use grid search to get the performance and using interpolation to get the smooth landscapes, we further provide some examples of the original performance heat-map in grids, as shown in Fig. 6
Figure 6: Metric landscapes in grid for CD.
@@
Appendix E Experimental Details and Additional Results
E.1 Experimental Details
E.1.1 Training Setting
Since LCSC employs weight checkpoints, we endeavored to replicate the training processes of the baseline models.
For DM, we follow DDIM (Song et al., 2020a) for the evaluation on CIFAR10 (Krizhevsky et al., 2009) and iDDPM (Nichol & Dhariwal, 2021) for the evaluation on ImageNet-64 (Deng et al., 2009). To improve the inference efficiency, we adopt DPM-Solver (Lu et al., 2022). For CM (Song et al., 2023), we evaluate LCSC with both CD and CT on CIFAR-10 and ImageNet-64, and CT on LSUN datasets. For CIFAR-10 and ImageNet-64, We train models with our own implementation. We follow all the settings reported by the official paper except ImageNet-64, on which we decrease the batch size to 256 on CM and 512 on DM due to the limited resources. We apply LCSC at different training stages. For LSUN datasets, we fine-tune the released official models and apply LCSC. We further train a LCM-LoRA model following the official setting and use LCSC with the final checkpoints. Specifically, for any selected iteration, we utilize checkpoints from every predetermined interval of iterations within a defined window size. Then we run a search process to find the optimal combination coefficients.
@@ E.1.1 Training Setting
For DMs, we were able to reproduce the results reported in the original papers successfully. However, for CMs, our training outcomes on CIFAR10 were slightly inferior to those documented in the original papers. Additionally, due to resource constraints, we opted for a smaller batch size than the original configuration when training on ImageNet. In the results tables, we present both our training outcomes and the results reported in the original papers, with the latter indicated as released. The specifics of the experimental setups are detailed below.
@@
For the experiments with DM, we utilize the DDIM (Song et al., 2020a) codebase (https://github.com/ermongroup/ddim) on CIFAR-10, adhering to the default configuration settings. For ImageNet-64, we employ the iDDPM (Nichol & Dhariwal, 2021) codebase (https://github.com/openai/improved-diffusion), setting the batch size to 512, the noise schedule to cosine, and maintaining other hyperparameters at their default values. During sampling, the DPM-Solver (Lu et al., 2022) (https://github.com/LuChengTHU/dpm-solver) is applied, conducting 15 timesteps of denoising using the default configuration for the respective setting. The reproduced results are on par with or slightly surpass those achieved using Euler integration, as reported in the original papers.
@@
For the experiments involving CM, we adhere closely to the configurations detailed in Table 3 of the original CM paper (Song et al., 2023). On CIFAR-10, instead of using the NCSNPP model (Song et al., 2020b) from CM’s official implementation, we opt for the EDM (Karras et al., 2022) architecture. This decision stems from the official CM code for CIFAR-10 being implemented in JAX while our experiments are conducted using PyTorch. For further details, refer to https://github.com/openai/consistency_models_cifar10. This difference in implementation may partly explain the discrepancies between our replication results and those reported in Table 1 of CM’s original paper (Song et al., 2023). On ImageNet-64 and LSUN, we follow the official implementation, employing the ADM (Dhariwal & Nichol, 2021) architecture. Due to resource constraints, we train the model with a smaller batch size of 256 instead of 2048 as used in the original study, resulting in worse outcomes. However, as indicated in Tab. 2, applying LCSC to models trained with a reduced batch size can still achieve performance comparable to models trained with a larger batch size.
E.1.2 Search Setting
@@
At each selected training iteration, historical weights are leveraged within designated window sizes and intervals. For CM, checkpoints have a window size of 40K with an interval of 100 on CIFAR-10, and a window size of 20K with an interval of 100 on ImageNet-64 and LSUN datasets. DM is assigned a window size of 50K and an interval of 200 for both datasets. FID calculation for DM on ImageNet-64 and CM on LSUN datasets utilizes a sample of 5K images, while 10K images are used for all other configurations. An evolutionary search spanning 2K iterations is applied consistently across all experimental setups.
E.1.3 Text-to-image Task
For text-to-image task, we fine-tune a LoRA based on the Stable Diffusion v1-5 model (Rombach et al., 2022) on CC12M dataset. We use a batch size of 12 and train the LoRA for 6k steps. For LCSC, we use the checkpoints saved between 4k and 6k steps with an interval of 20. When sampling, we insert LoRA to Dreamshaper-7, which is a fine-tuned version of Stable Diffusion v1-5, and use 1k samples from CC12M to calculate FID. Finally, we randomly sample 10k samples from MS-COCO dataset to evaluate the search results.
For baselines, we find that the performance is sensitive to the scale of LoRA. Therefore, we first conduct a coarse scan to determine the approximate range of the optimal LoRA scale, followed by a fine sweep. We find 0.15 is the optimal scale for LoRA and use it as the baseline.
E.2 Two-phase Convergence
Fig. 8 shows that the generation quality convergence of DM and CM can be divided into two phases. The initial phase is relatively brief, during which DM and CM rapidly acquire the capability to generate visually satisfactory images. In contrast, the second phase is characterized by a slower optimization of models, focusing on the enhancement of sample quality. It is noteworthy that the majority of training iterations belongs to the second phase.
E.3 Search Cost Estimation
@@
Next, we examine the overall training and search time costs as outlined in Tab. 9, and we delve into the speedup ratio detailed in Tabs. 10 and 11. Given that DMs necessitate multiple timesteps for sampling, the overall search cost for DM proves to be non-negligible when compared to the training cost. This results in no observable speedup in convergence upon applying LCSC. However, this search cost is considered manageable and warranted for the anticipated improvements to the final model convergence.
In contrast, for CMs, the search cost is substantially lower than the training cost, which often results in a markedly faster convergence when combining training with LCSC than with training alone. The precise speedup ratios for CM training are detailed in Tabs. 10 and 11. It should be noted that the figures in these tables may exhibit slight discrepancies from those reported in Sec. 5 due to the latter being profiled in an environment with more variables and potential fluctuations.
@@
E.4 Hyperparameter Study
The hyperparameters of LCSC includes the window size ($n_{K}-n_{1}$) for retrieving the historical weights, the interval ($n_{k}-n_{k-1}$) for two adjacent weight checkpoints, the number of samples for computing FID and the number of search iterations, and the number of timesteps for DM.
Tab. 12 illustrates the impact of varying the number of samples, search iterations, and the size of the interval. We observe that increasing the number of samples and search iterations leads to higher performance, though this comes at the expense of increased search cost. In practice, we find that 10K samples and 2K search iterations can efficiently identify strong models. With a fixed window size, the interval between two checkpoints determines the dimension of the search space. Our findings suggest that search performance generally improves as the search space expands, but limiting the search dimension to fewer than 200 can detrimentally affect search performance.
(a) Results on DM
Samples Search Iters. Interval FID($\downarrow$) IS($\uparrow$)
10K 2K 200 3.87 9.27
2K 2K 200 9.23 9.11
5K 2K 200 4.04 9.09
10K 1K 200 4.01 9.17
10K 4K 200 3.79 9.28
10K 2K 100 3.82 9.15
10K 2K 500 4.41 9.11
(b) Results on CM
Samples Search Iters. Interval FID($\downarrow$) IS($\uparrow$)
10K 2K 100 2.76 9.71
2K 2K 100 3.40 9.39
5K 2K 100 2.79 9.66
10K 1K 100 3.03 9.57
10K 4K 100 2.69 9.70
10K 2K 200 2.84 9.71
10K 2K 50 2.71 9.79
For the impact of window size $(n_{K}-n_{1})$ and sampling NFE of DM, the findings are detailed in Tab. 13. The results reveal that, despite earlier models being further from convergence, a sufficiently large window for accessing historical weights proves advantageous. Furthermore, conducting a lower NFE during the search in the DM context results in a similar reduction in FID but a smaller improvement in IS. This suggests that the search output’s generality across different metrics diminishes when the generated samples during the search are less accurate, i.e., exhibit significant truncation error.
@@ E.4 Hyperparameter Study
E.5 Detailed Results of EMA Rate Grid Search
For the final training model of each experimental configuration, we conduct a comprehensive sweep across a broad range of EMA rates, presenting the optimal outcomes in Sec. 5. The exhaustive results are detailed in Tab. 14. In the majority of scenarios, the default EMA rate employed in the official implementations of our baseline models (Song et al., 2023; Ho et al., 2020; Nichol & Dhariwal, 2021) yields slightly inferior performance compared to the best EMA rate identified. Nevertheless, the findings demonstrate that LCSC consistently surpasses all explored EMA rates, underscoring the limitations of EMA as a strategy for weight averaging.
@@ E.5 Detailed Results of EMA Rate Grid Search
E.6 Different formulation of the base checkpoints
As discussed in Sec. 4.2, we subtract the first checkpoint from each checkpoint and use these differences as the weighted base. In this section, we explore an alternative approach for defining the base checkpoints: using all the checkpoints themselves as the base. In this case, the final combined weight is:
$\bm{\alpha}\cdot\Theta=\sum^{K}_{i=1}\alpha_{i}\bm{\theta}_{n_{i}}.$ (77)
We apply this formulation in our search, with the results presented in Tab. 15 and Tab. 16. We can see that the difference search method outperforms the checkpoint search method in all cases on the ImageNet-64. On CIFAR-10, the performance of the two methods are similar with each other.
Table 15: The comparison of different formulations for base checkpoints on CIFAR-10 dataset. LCSC-Diff stands for the formulation we use in our main experiments. LCSC-Direct stands for using the checkpoints themselves as the weighted base.
CD LCSC-Direct 50K 512 1 3.18 9.60 0.67 0.58 $\sim$14$\times$
250K 512 1 2.76 9.71 0.67 0.59 $\sim$3.1$\times$
800K 512 1 2.42 9.76 0.67 0.60 -
800+40K 512 1 2.38 9.70 0.67 0.60 -
100K 128 1 3.34 9.51 0.67 0.57 $\sim$23$\times$
LCSC-Diff 50K 512 1 3.10 9.50 0.66 0.58 $\sim$14$\times$
250K 512 1 2.66 9.64 0.67 0.59 $\sim$3.1$\times$
800K 512 1 2.44 9.82 0.67 0.60 -
800+40K 512 1 2.50 9.70 0.68 0.59 -
100K 128 1 3.21 9.48 0.66 0.58 $\sim$23$\times$
CT LCSC-Direct 400K 512 1 8.60 8.89 0.67 0.47 $\sim$1.9$\times$
800+40K 512 1 8.05 8.98 0.70 0.45 -
450K 128 1 8.33 8.67 0.69 0.44 $\sim$7$\times$
LCSC-Diff 400K 512 1 8.89 8.79 0.67 0.47 $\sim$1.9$\times$
800+40K 512 1 7.05 9.01 0.70 0.45 -
450K 128 1 8.54 8.66 0.69 0.44 $\sim$7$\times$
Table 16: The comparison of different formulations for base checkpoints on ImageNet-64 dataset. LCSC-Diff stands for the formulation we use in our main experiments. LCSC-Direct stands for using the checkpoints themselves as the weighted base.
CD LCSC-Direct 300K 256 1 5.71 41.8 0.68 0.62 $\sim$15$\times$
600+20K 256 1 5.54 40.9 0.68 0.62 $\sim$7.6$\times$
LCSC-Diff 300K 256 1 5.51 39.8 0.68 0.62 $\sim$15$\times$
600+20K 256 1 5.07 42.5 0.69 0.62 $\sim$7.6$\times$
CT LCSC-Direct 600K 256 1 12.1 35.1 0.67 0.54 $\sim$10.4$\times$
800K 256 1 11.1 35.7 0.65 0.57 $\sim$6.3$\times$
LCSC-Diff 600K 256 1 10.5 36.8 0.66 0.56 $\sim$10.4$\times$
800K 256 1 9.02 38.8 0.68 0.55 $\sim$7.3$\times$
E.7 Evaluation with Other Metrics
E.7.1 Evaluation with FCD and KID
In this section, we present the results of FCD and KID using CT method on CIFAR-10 and ImageNet-64 dataset. The results are shown in Tab. 17.
Table 17: Evaluation results with FCD and KID metrics on CIFAR-10 and ImageNet-64 datasets.
(a) Results on CIFAR-10.
Method Training Iter Batch Size FID($\downarrow$) FCD($\downarrow$) KID($\downarrow$)
CT 400k 512 12.1 43.0 7.47e-3
CT 800k 512 9.87 35.8 5.14e-3
LCSC 450k 128 8.54 27.2 4.65e-3
LCSC 400k 512 8.89 34.4 3.96e-3
LCSC 800+40k 512 7.05 24.9 2.90e-3
(b) Results on ImageNet-64.
Method Training Iter Batch Size FID($\downarrow$) FCD($\downarrow$) KID($\downarrow$)
CT 600k 256 16.6 49.8 9.22e-3
CT 800k 256 15.8 47.5 8.69e-3
CT 800k 2048 13.1 47.5 8.55e-3
LCSC 600k 256 10.5 38.8 4.41e-3
LCSC 800k 256 9.02 30.0 3.87e-3
E.7.2 Evaluation with FID on Test Dataset
In this section, we report the FID calculated on the test dataset to validate that LCSC does not over-fit on the training data. Results are shown in Tab. 18. We can see that LCSC also achieves significant improvement on test FID, indicating its generalization ability across different data.
Table 18: Evaluation results on CIFAR-10 test set.
Method Training Iter Batch Size FID($\downarrow$) test FID($\downarrow$)
CD 800k 512 3.66 5.88
CD 840k 512 3.65 5.85
LCSC 100k 128 3.21 5.48
LCSC 250k 512 2.66 4.88
LCSC 800k 512 2.44 4.70
LCSC 800+40k 512 2.50 4.75
E.8 Search Results with Stable Diffusion Models
To further demonstrate the effectiveness of LCSC in enhancing DMs, we conduct experiments on Stable Diffusion (Rombach et al., 2022) checkpoints. Specifically, we fine-tune the Stable Diffusion v1-5 model on CC12M using LoRA for 20k iterations. We then apply LCSC to the saved checkpoints at intervals of 100 iterations. The results, presented in Tab. 19, show that LCSC achieves a significant improvement compared to the released Stable Diffusion checkpoints with the same sampling NFE.
To investigate whether LCSC can accelerate the Stable Diffusion model, we further test PickScore (Kirstain et al., 2023) between the LCSC model with 10 NFE and the Stable Diffusion model with 15 NFE. The results are 0.49 and 0.51 for Stable Diffusion and LCSC, respectively, with a 57% winning rate for LCSC. This further demonstrate the ability of LCSC to accelerate inference speed for diffusion models.
Table 19: Results of LCSC on Stable Diffusion Checkpoints.
NFE FID PKS WR@PKS CLIP Score
LCSC 15 16.30 0.53(v.s. SD)/0.53(v.s. LoRA) 55%(v.s. SD)/56%(v.s. LoRA) 26.69
SDv1-5 15 17.55 0.47 44% 26.60
LoRA tuning 15 17.05 0.47 45% 26.61
LCSC 10 16.68 0.59(v.s. SD)/0.51(v.s. LoRA) 64%(v.s. SD)/53%(v.s. LoRA) 26.61
SDv1-5 10 18.16 0.41 36% 26.57
LoRA tuning 10 17.35 0.49 47% 26.56
Appendix F More Insights and Analysis
F.1 Analysis of Search Patterns
F.1.1 More Examples of Search Patterns
@@
We demonstrate search results on ImageNet-64 in Fig. 9, which share similar patterns with results on CIFAR-10 (refer to Fig. 5). Additionally, Fig. 10 shows that the checkpoints assigned with larger coefficients often has lower FID than the checkpoints with small coefficients.
@@
Figure 10: FID of checkpoints with varying magnitude of coefficients.
F.1.2 Search Pattern of different Random Seed
We compare search patterns across different random seeds. As shown in Fig. 11, although each seed’s dominant coefficients map to distinct subsets of weight checkpoints, they produce similar outcomes in terms of FID as provided in the caption. These results suggest multiple high-quality basins exist within the weight checkpoint subspace, and LCSC converges to one of them depending on the randomness. Previous study demonstrates that local minimal in networks are connected by simple curves over which losses are nearly constant (Garipov et al., 2018). To investigate whether solutions found by LCSC are also connected we average the search patterns obtained from multiple random seeds and evaluate the performance of the resulting model. As shown in Fig. 12, the averaged search pattern becomes homogeneous and contains smaller coefficients, yet it achieves comparable performance, as detailed in the figure caption. This observation suggests that the search results of LCSC may also be connected by low-loss curve.
Figure 11: Visualization of linear combination coefficients obtained by LCSC with different random seeds using CD on CIFAR10. Their FID scores are: 2.76 (seed 1) and 2.69 (seed 2).
Figure 12: Visualization of averaged coefficients obtained by LCSC across different random seeds using CD on CIFAR10. Their FID scores are: 2.76 (left), 2.70 (middle), 2.83 (right).
F.1.3 The Critical Role of Search Space Configuration
As explored in Sec. 6, restricting averaged weights to the convex hull of candidate checkpoints, where all combination coefficients are non-negative and their sum equals 1, might limit search efficacy. To test this hypothesis, we perform searches under convex combination conditions by clipping coefficients to non-negative values and normalizing their sum to 1. Tab. 20 presents these findings, with “w/o” indicating no restriction on coefficients (allowing values below zero) and “w” representing the convex hull restriction. The results clearly show superior performance without the convex restriction, underscoring its limiting effect on search outcomes. Furthermore, we discover that normalizing the coefficient sum to 1 is vital for effective search exploration, a practice we continue even after lifting the convex condition, as detailed in Eq. 7.